### v1

# Enhancing Negation Awareness in Universal Text Embeddings: A Data-efficient and Computational-efficient Approach

Abstract
Negation plays an important role in various natural language processing tasks such as Natural Language Inference and Sentiment Analysis tasks. Numerous prior studies have found that contextual text embedding models such as BERT, ELMO, RoBERTa or XLNet face challenges in accurately understanding negation. Recent advancements in universal text embeddings have demonstrated superior performance over contextual text embeddings in various tasks. However, due to the bias in popular evaluation benchmarks, the negation awareness capacity of these models remains unclear. To bridge the gap in existing literature, an in-depth analysis is initiated in this work to study the negation awareness of cutting-edge universal text embedding models. Our findings reveal a significant lack of negation awareness in these models, often interpreting negated text pairs as semantically similar. To efficiently deal with the conflict that different tasks need different trade-offs between topic and negation information among other semantic information, a data-efficient and computational-efficient embedding re-weighting method is proposed without modifying the parameters of text embedding models. The proposed solution is able to improve text embedding models’ negation awareness significantly on both simple negation understanding task and complex negation understanding task. Furthermore, the proposed solution can also significantly improve the negation awareness of Large Language Model based task-specific high dimensional universal text embeddings.
Index Terms:
Negation awareness, Semantic similarity, Universal text embeddings, Natural Language Processing
I Introduction
Text embedding has gained significant attention from both industry and academia due to its crucial role in various Natural Language Processing (NLP) tasks such as information retrieval [1], sentiment analysis [2, 3], text classification [4], text clustering [5, 6], etc. Recently, the importance of text embeddings has been further highlighted in the context of Large Language Models (LLMs) based applications, such as Retrieval Augmented Generation (RAG) systems. This is primarily because these applications rely on high-quality text embeddings for performing vector search, which involves retrieving the most relevant contexts/documents for LLM Question Answering (QA) tasks [7].
One crucial aspect related to the effectiveness of text embedding model is the comprehension and interpretation of negation information, a key linguistic construct that significantly influences sentiment and meaning [8]. Its accurate interpretation is therefore critical in sentiment analysis and Natural Language Inference (NLI) tasks [9, 10]. However, several previous works have shown that contextual text embeddings like BERT [11] lack understanding of negations and fail to attribute sufficient importance to the word “not” [12, 13, 10]. Such bias could render the entire output of the model invalid, thereby diminishing the reliability of such systems, particularly in scenarios involving opinion mining, fact-checking or measuring answer correctness [8].
The field of universal text embeddings has recently witnessed remarkable advancements. This progress can be attributed to the advancements in the quantity and quality of diverse text datasets across different tasks [14, 15], synthetic data generated by different LLMs [16, 17], using LLMs as backbones and the emergence of benchmarks with the focus on novel task and domain generalization such as the Massive Text Embedding Benchmark (MTEB) [18]. Representative universal embedding models such as GTE [19], BGE [14], E5 [20, 17, 21], Gecko [16] or LLM2Vec [22] can effectively address diverse downstream tasks without fine-tuning. They have demonstrated significant improvements over contextual text embeddings across diverse tasks including information retrieval, reranking, clustering and pair classification tasks [7]. However, evaluation benchmarks such as MTEB are biased towards tasks such as retrieval, classification and clustering [18, 7, 23]. The negation awareness of state-of-the-art universal text embeddings remains unclear.
This work bridges the gap in existing literature and makes the following contributions:
  • Novel in-depth analysis is proposed to understand how similar are negated text pairs interpreted by diverse state-of-the-art universal text embeddings.
  • To deal with the conflict that different tasks need different trade-offs between topic and negation information among other semantic information, a data-efficient and sustainable embedding re-weighting method is proposed without fine-tuning.
  • We demonstrate that the proposed solution is able to improve text embedding models’ negation awareness significantly on both simple negation understanding task and complex negation understanding task.
  • We further demonstrate that the proposed solution can also significantly improve the negation awareness of LLM-based task-specific high dimensional universal text embeddings.
II Related Works
II-A Universal text embeddings
The field of text embeddings has seen substantial improvements in recent years. Text embedding models can be categorized in four eras [7]: the first era stands for count-based embeddings such as Bag of Words [24] and Term Frequency-Inverse Document Frequency (TF-IDF) [25] which are transformed into low-dimensional dense embeddings with methods like Latent Semantic Indexing (LSA) [26]; the second era is static dense word embeddings represented by Word2Vec [27], GloVe [28] and FastText [29] which can not deal with polysemy; the third era is contextualized embeddings where context-sensitive dynamic embeddings can adapt or change based on context and deal with polysemy, represented by Embeddings From Language Models (ELMo) [30], Generative Pre-trained Transformer (GPT) [31] and Bidirectional Encoder Representations from Transformers (BERT) [11]. Recently, there has been significant progress in the fourth generation of universal text embeddings which aim at building an unified comprehensive text embedding to address a multitude of input text length, downstream tasks, domains and languages [7].
To achieve universality, large quantity datasets with diverse mixtures are used for both pre-training and fine-tuning by GTE [19]. BGE [14] focuses on improving the training data quality by filtering irrelevant text pairs and using multi-task high quality labeled data for fine-tuning. Similarly, E5 [20] also focuses on data quality improvement by constructing a curated web-scale text pair dataset named Colossal Clean text Pairs (CCPairs), while Multilingual E5 [21] combines diverse real world datasets with synthetic datasets containing 150k unique instructions and 93 languages generated by GPT-3.5/4 in order go generalize across different tasks as well as different languages. Some other universal text embeddings focus on new loss functions including Universal AnglE Embedding (UAE) [32] and mxbai-embed-large-v1 from [33, 34]. LLMs are also used as backbones for universal text embeddings (e.g. E5-Mistral [17], LLM2Vec [22], gte-Qwen1.5-7B-instruct [19]) as they do not need the large-scale contrastive pre-training.
II-B Negation understanding
Figure 1: Example of semantic similarity based on universal text embeddings. S1 and S2 are paraphrases from STSB [35] with human annotated similarity of 0.96. S3 is a simple negation of the reference S1. The cosine similarity between text embeddings of S1 and S2 as well as S1 and S3 are compared.
Negation plays an important role in many NLP tasks including NLI [8], sentiment analysis [9], Natural Language Understanding (NLU) tasks [10], etc. The complexity of negation, as highlighted by [9], presents substantial challenges for commercial sentiment analysis. Contextual text embedding models such as BERT [11] is shown to be insensitive to negations in [12] through a completion task involving masked hypernyms in affirmative and negated versions of each sentence. Other Contextual text embeddings including Transformer-XL [36] and ELMO [30] are also found to exhibit poor performance in distinguishing between affirmative and negative sentences when subjected to negative cloze tests [37]. Moreover, they tend to make identical predictions regardless of the polarity of the sentence [38]. In the study conducted by [13, 10], novel benchmark was created to evaluate the ability of contextual embedding models in terms of recognizing negation. The experimental findings revealed that even after fine-tuning with additional negation-containing pairs, contextual text embeddings (RoBERTa [39], XLNet [40] and BERT [11]) still face challenges in accurately understanding negation.
Several studies have been conducted in order to improve the negation awareness of contextual text embedding models such as BERT. In [41], BERTNOT is proposed by fine-tuning a pre-trained BERT model using a knowledge distillation objective as well as a novel unlikelihood objective, which is based on negated generic sentences extracted from a raw text corpus. The proposed BERTNOT has been observed to effectively incorporate negation knowledge into the BERT-base model. However, it has been found to be ineffective with the BERT-large model. A novel negation-focused pre-training strategy is proposed in [42] to improve the performance and generalization ability of negation detection using targeted data augmentation and negation masking. Another new pre-training framework with a weighted cross-entropy loss and elastic weight consolidation regularization was proposed in [43] to overcome catastrophic forgetting and augment the negation knowledge in contextual text embedding models.
Compared to contextual embeddings, universal embeddings have the capacity of generalizing across diverse tasks. However, their negation awareness is yet to be determined as the evaluation benchmarks such as MTEB are biased towards tasks with many datasets, notably retrieval, classification and clustering [18, 7]. A simple example is shown in Figure 1: given the reference sentence S1 (’Polar bear DNA may help fight obesity’) and a semantic similar sentence S2 (’Polar bear study may boost fight against obesity’) from Semantic Textual Similarity Benchmark (STSB) [35] with the groudtruth similarity value 0.96, S3 is the negated version of S1. Diverse universal text embeddings are used to measure the semantic similarity between S1 & S2 and the semantic similarity between S1 & S3. While the semantic similarity between S1 & S2 is expected to be much larger than the one between S1 & S3, most universal text embeddings fail with this simple example. To understand better the negation awareness of the state-of-the-art universal text embeddings, holistic analysis and comparisons are conducted in the following sections.
III How similar are negated text pairs for universal text embeddings?
Semantic similarity refers to the degree of overlap in meaning [35]. Measuring the semantic similarity using cosine over text embeddings is one of the most adopted method. Among the 8 evaluation tasks from MTEB [18], embedding based semantic similarity is used by 7 of them. Recently, text embedding based semantic similarity is also used as evaluation metrics to measure the answer correctness of RAGs or LLMs [44, 45]. Despite its wide usage, there is few study trying to understand what kind of semantic information is encoded and highlighted in the semantic similarity based on text embeddings. In the literature of Semantic Textual Similarity (STS) studies such as [46, 35], semantic similarity is defined as (re-scaled to the range between 0 and 1):
  • 0 means the the pair of texts are on different topics;
  • 0.2 means the the pair of texts are not equivalent, but are on the same topic;
  • 0.4 means the the pair of texts are not equivalent, but share some details;
  • 0.6 means the the pair of texts are roughly equivalent, but some important information differs/missing;
  • 0.8 means the the pair of texts are mostly equivalent, but some unimportant details differ;
  • 1 means the the pair of texts are completely equivalent;
One limitation of this definition is that: the semantic similarity is mainly based on topic information while other semantic information such as negation is not included. When it comes to two pieces of text where one is the negated version of the other, the situation becomes more complex. For instance, consider these two sentences: (1) ”The horse is white.” and (2) ”The horse is not white.”: from a purely syntactic perspective, the sentences are very similar. They share the same subject and predicate, with the only difference being the presence of negation in (2). From a purely topic perspective, the sentences are similar as they share similar topics: the color of the horse. However, semantic textual similarity is not just about the structure or the words used, but about the meaning conveyed by the sentences. The negation in the second sentence changes the meaning entirely. According to [43], the significance of negation in understanding the semantic relationship in natural language texts is profound, as it reverses the overall semantic meaning of a sentence. Consequently, the primary focus of this section is to study how universal text embeddings interpret the semantic similarity between negated text pairs.
III-A Dataset
STSB [35] is a collection of English datasets that have been used in the SemEval and *SEM STS shared tasks over the period from 2012 to 2017 [35]. Annotation of text pair’s similarity is accomplished via crowdsourcing, integrating pragmatic and global knowledge, thereby enhancing their interpretability and utility for subsequent tasks [35]. Within the scope of this study, the human annotation similarity scores in the STSB dataset, which typically range from 0 to 5, are re-scaled to the range between 0 and 1. Given STSB text pairs with the human annotated semantic similarity in the form of [$sentence\_1_{i}$, $sentence\_2_{i}$, $similarity\_score_{i}$], a sentence-level negation tool (focuses on verbal negations and supports the simple addition and deletion of negation cues) developed by [8] is used to create $neg\_sentence\_1_{i}$: the simple negated version of $sentence\_1_{i}$ (e.g. transform ”The horse is white” into its simple negated version of ”The horse is not white”). The training partition of the dataset is used for analysis (in this section) and training (in next section), while the development and test sets are combined for testing (in next section).
III-B Universal text embeddings
In this study, diverse universal text embeddings with different parameters size, training data, loss functions and fine-tuning strategies are selected. The focus has been on those that are open-source and have demonstrated superior performance on the MTEB benchmark, including: gte-large-en-v1.5 (434M) and gte-base-en-v1.5 (137M) [19], bge-large-en-v1.5 (335M) and bge-base-en-v1.5 (109M) [14], e5-large-v2 (335M) and e5-base-v2 with (109M) [20], multilingual-e5-large-instruct [21] (560M) and mxbai-embed-large-v1 [33, 34] (335M). Two widely used contextual text embeddings from SentenceTransformers [6] are selected as baseline models, including all-mpnet-base-v2 (110M) and all-roBERTa-large-v1 (355M).
III-C Experiments
Figure 2: The histogram plot of semantic similarity scores based on universal text embeddings (columns): row 1 to row 5 represents the results of group 1 to group 5. In each subplot, the x-axis is the similarity values between 0 and 1, while y axis is the sample size (in percentage). Blue color represents the semantic similarity between $sentence\_1_{i}$ and $sentence\_2_{i}$; orange color represents the semantic similarity between $sentence\_1_{i}$ and $neg\_sentence\_1_{i}$.
In order to study how universal text embeddings interpret negation, the training data of STSB are separated into five groups based on the human annotated similarity between $sentence\_1_{i}$ and $sentence\_2_{i}$: group 1 (0-0.2), group 2 (0.2-0.4), group 3 (0.4-0.6), group 4 (0.6-0.8), and group 5 (0.8-1). The semantic similarity $Sim\_12$ between S1 ($sentence\_1$) and S2 ($sentence\_2$) as well as the semantic similarity $Sim\_1neg1$ between S1 and its negated version NegS1 ($neg\_sentence\_1$) within each group based on each universal text embeddings are calculated. By comparing $Sim\_1neg1$ with $Sim\_12$ relatively within each group, we can get the insights on how similar are negated text pairs interpreted by the state-of-the-art universal text embeddings.
The histogram of $Sim\_12$ within each group is illustrated in blue color in Figure 2: row 1 to row 5 represents the results from group 1 to group 5 and each column represents a text embedding model. Generally speaking, the absolute similarity values do not match the human annotation. For instance, the human annotation for group 1 is between 0 and 0.2, while the similarity values calculated by most embedding models are between 0.2 and 0.8. For E5 series, similarity values are between 0.6 and 1 for group 1. However, the evaluation metric used by STSB is Pearson correlation between embedding based semantic similarity scores and human annotations [35]. Hence, the relative position should also be compared. For all universal text embeddings, the histogram of $Sim\_12$ moves towards the right (more similar) from data in group 1 (dissimilar text pairs) to data in group 5 (similar text pairs). This result indicates that the universal text embeddings based semantic similarity is correlated with the human annotations.
The histogram of $Sim\_1neg1$ within each group is illustrated as orange color in Figure 2: the semantic similarity values are within the range [0.6, 1] for all groups and all universal text embeddings. In order to answer how similar are negated text pairs interpreted by the state-of-the-art universal text embeddings, the histogram of $Sim\_1neg1$ is compared with $Sim\_12$ within in each group. For Group 1: The human annotated similarity between S1 and S2 is within [0, 0.2], which means they are very dissimilar. From the first row of Figure 2, it can be seen that S1 is more similar to its negated version NegS1 (orange color) than to S2 (blue color) across all universal text embeddings (in 99.27% cases on average). Similar observations can be found for Group 2 and 3. For Group 4: The human annotated similarity between S1 and S2 is within [0.6, 0.8] which means they are roughly similar. From the fourth row of Figure 2, it can be seen that the similarity distribution of $Sim\_1neg1$ overlaps with $Sim\_12$. On average (across all embedding models), S1 is more similar to NegS1 than S2 in 51.57% cases. In Group 5: The human annotated similarity between S1 and S2 is within [0.8, 1] which means they are very similar. From the last row of Figure 2, it can be seen that S1 is more similar to S2 than NegS1 (in 77.25% cases). These observations reveal that the distribution of $Sim\_12$ and $Sim\_1neg1$ overlaps most within Group 4, suggesting that these embeddings lack negation awareness, interpreting negated text pairs as roughly similar in terms of semantic meaning. This aligns with the definition of human-annotated semantic similarity scores, as Group 4 has scores ranging from 0.6 to 0.8, indicating that the negated and original sentences are equivalent in terms of topic and syntax, but differ in the presence of negation. This suggests that topic and syntactic information are prioritized over negation information in defining semantic similarity for these embeddings. However, for NLI or sentiment analysis tasks, understanding negation is crucial as it can reverse sentiment entirely. Therefore, adjusting the trade-off among different semantic information such as topic and negation presents a significant challenge for universal text embeddings in achieving task universality.
IV Improve negation awareness in universal text embeddings
From the analysis in the previous section, it can be concluded that the negation awareness of the state-of-the-art universal text embeddings is limited. Previous studies [41, 42, 43] assume that the embedding models lack of negation knowledge and focus on improving negation awareness by fine-tuning pre-trained models on negation datasets. However, fine-tuning on negation datasets may hurt the performance of universal text embeddings on other tasks (e.g. retrieval or topic modeling) which requires a better understanding of topic information. To efficiently deal with such conflicts, a data-efficient and sustainable embedding re-weighting function is proposed in this section without modifying the parameters of text embedding models. The main hypothesis is that universal text embeddings inherently encapsulate a broad but imbalanced spectrum of semantic information including topics, sentiment, negation and so on due to their large scale training/fine-tuning on extensive and diverse datasets. By assigning greater weights to the dimensions that predominantly capture negation-related information, they are expected to recognize and process negations more effectively. To ensure the universality across different tasks: this negation-aware weight will only be applied for negation related tasks, while the original unweighted embedding is used for other tasks.
Given the D-dimensional embedding $E_{T_{i}}=\{e_{T_{i}}^{0},...e_{T_{i}}^{D}\}$ of text $T_{i}$, $E_{P_{i}}$ is the embedding of $P_{i}$ (the paraphrased version of $T_{i}$), $E_{N_{i}}$ is the embedding of $N_{i}$ (the negated version of $T_{i}$). The cosine similarity between $T_{i}$ and $P_{i}$ is:
$\begin{split}Cos(T_{i},P_{i})&=\frac{e_{T_{i}}^{1}\times e_{P_{i}}^{1}+...+e_{T_{i}}^{k}\times e_{P_{i}}^{k}+...+e_{T_{i}}^{D}\times e_{P_{i}}^{D}}{||E_{T_{i}}||\times||E_{P_{i}}||}\\ &=u_{(T_{i},P_{i})}^{1}+...+u_{(T_{i},P_{i})}^{k}+...+u_{(T_{i},P_{i})}^{D}\\ &where:\quad u_{(T_{i},P_{i})}^{k}=\frac{e_{T_{i}}^{k}\times e_{P_{i}}^{k}}{||E_{T_{i}}||\times||E_{P_{i}}||}\end{split}$ (1)
Given the triplet $[T_{i},P_{i},N_{i}]$, ideally we want the similarity between the paraphrased text pair $Cos(T_{i},P_{i})$ to be greater than the similarity between the negated text pair $Cos(T_{i},N_{i})$:
${Diff}_{i}=Cos(T_{i},P_{i})-Cos(T_{i},N_{i})=\Sigma_{k=1}^{D}v_{i}^{k}>0$ (2)
where: $v_{i}^{k}=u_{(T_{i},P_{i})}^{k}-u_{(T_{i},N_{i})}^{k}$. If $v_{i}^{k}$ is larger, the dimension $k$ contributes more to the negation awareness, hence the dimension $k$ should be assigned with a larger weight. Given training data with $N$ triplets, the mean contribution of each dimension $k$ can be calculated as:
$\overline{v}^{k}=\frac{1}{N}\times\Sigma_{i=1}^{N}v_{i}^{k}$ (3)
Given the average contribution of each dimension $\overline{V}=[\overline{v}^{1},...\overline{v}^{D}]$ (re-scaled by dividing $max(\overline{V})$ ), the proposed weighting function is:
$w^{k}=\frac{e^{a\times\overline{v}^{k}}}{\Sigma_{i=1}^{D}e^{a\times\overline{v}^{i}}}$ (4)
where $a$ is a hyperparameter to control the trade off between negation information and general semantic information in text embeddings. One toy example of the the average contribution of each dimension to negation awareness $\overline{V}$ is shown in Figure 3 (a), the corresponding dimension weights when applying Equation 4 are shown in Figure 3 (b), (c), (d) when $a$ equals to 2, 4, 6 respectively. It can be seen that larger $a$ assigns larger weights on to the dimension with positive contribution to negation awareness.
For negation related tasks, the dimension weights are then applied through element-wise multiplication to the original embeddings. This is equivalent to apply a weighted cosine similarity on the original embeddings.
Figure 3: A toy example of $\overline{V}$ with 100 dimensions is shown in (a): X-axis is the embedding dimension, Y-axis is the corresponding contribution to negation awareness. The impact of hyperparameter $a$ on the dimension weights are shown in (b), (c), (d) when $a$ equals to 2, 4, 6 correspondingly: X-axis is the dimension, Y-axis is the corresponding weight.
V Experimental results
V-A Experiments on simple negation
V-A1 Datasets
In order to test the effectiveness of the proposed embedding re-weighting method, a paraphrase detection task is formulated using the STSB dataset: the Group 5 data from STSB [$sentence\_1_{i}$, $sentence\_2_{i}$] are used as paraphrase text pairs. The embedding models are then required to identify the accurate paraphrase of $sentence\_1_{i}$ from [$sentence\_2_{i}$, $neg\_sentence\_1_{i}$]. The training split of STSB is used as training data, the validation split and test split of STSB are combined as test data in this work.
V-A2 Experiment Setups
The training data is used to select the best hyperparameter $a$ with grid search. The search space of hyperparameter $a$ in this work ranges from 0 to 5, incrementing in steps of 0.25. Two evaluation metrics are used in this experiment:
  • Accuracy: measures if state-of-the-art text embeddings can correctly identify semantically similar text $sentence\_2_{i}$ from negated text $neg\_sentence\_1_{i}$ which has similar surface form as $sentence\_1_{i}$.
  • Correlation: measures if the similarity score of state-of-the-art text embeddings correlates well with human annotated similarity values using Pearson correlation following the evaluation protocol of STSB.
V-A3 Experimental results
TABLE I: Performances of text embeddings on STSB dataset: corr is the correlation between original embedding similarity and the human annotations, $corr\_W_{stsb}$ is the correlation between weighted embedding similarity and the human annotations; acc is the accuracy of original embedding; $acc\_W_{stsb}$ is the accuracy of weighted embedding.
models corr $corr\_W_{stsb}$ acc $acc\_W_{stsb}$ gte-base-en-v1.5 87.29 86.35 (-0.94) 79.27 80.18 (+0.91) gte-large-en-v1.5 84.49 83.41 (-1.08) 74.55 79.64 (+5.09) bge-base-en-v1.5 86.80 85.95 (-0.85) 75.09 79.82 (+4.73) bge-large-en-v1.5 87.72 86.67 (-1.05) 79.82 86.73 (+6.91) e5-base-v2 86.63 85.56 (-1.07) 72.55 78.00 (+5.45) e5-large-v2 86.26 85.58 (-0.68) 76.73 79.27 (+2.54) multilingual-e5 85.26 84.06 (-1.20) 76.36 82.55 (+6.19) mxbai-embed-large-v1 88.57 87.70 (-0.87) 83.64 89.27 (+5.63) Average (Universal) 86.63 85.66 (-0.97) 77.25 81.93 (+4.68) all-mpnet-base-v2 86.37 85.48 (-0.89) 71.45 74.18 (+2.73) all-roBERTa-large-v1 86.82 85.58 (-1.24) 73.82 78.73 (+4.91) Average (Contextual) 86.60 85.53 (-1.07) 72.63 76.46 (+3.83)
The experimental results of state-of-the-art text embeddings on STSB dataset are shown in Table I: acc is the test accuracy of original embedding; $acc\_W_{stsb}$ is the accuracy of weighted embedding where the weight is calculated following Equation 4. Generally speaking, universal text embeddings have better performance than the two baseline contextual text embedding models. The proposed dimension weighting method improves the accuracy across all text embeddings with an average improvement of 4.51%. bge-large-en-v1.5 benefits the most from the proposed method with an improvement of 6.91% while gte-base-en-v1.5 has the smallest improvement of 0.91%. To study if improved negation awareness would hurt the topic information understanding, the correlation between human annotated similarity values (mainly based on topic information) and embedding semantic similarity (before and after applying weights) is also calculated as shown in Table I: corr is the correlation between original embedding similarity and the human annotations, $corr\_W_{stsb}$ is the correlation between weighted embedding similarity and the human annotations. On average, the semantic similarity based on weighted embeddings have slightly reduced correlation (1% less) with human annotations compared to the original embedding.
In general, universal embeddings benefit more from the proposed solution (+4.68 improvement) than contextual embeddings (+3.83 improvement). In terms of embedding model parameter size: larger text embedding models generally have larger dimensions and better negation awareness than smaller embedding models, with the exception of GTE series: gte-base-en-v1.5 has better performance on both semantic textual similarity (corr) and negation awareness (acc) than gte-large-en-v1.5.
V-B Generalization ability of weights learnt from STSB to SemAntoNeg (more complex negation)
SemAntoNeg Benchmark: This benchmark from [38] contains more complex negation forms. It assesses the proficiency of embedding models in accurately depicting phrases that incorporate both antonymy and negation. Through a paraphrase task, the model must identify the correct paraphrase among three candidate expressions, each featuring negated sentences and antonym substitutions. Thus, embedding models must understand that adding or removing a negation, coupled with an antonym substitution, retains the sentence’s original semantic meaning. Unlike STSB, there is no human annotation of semantic similarity on SemAntoNeg. Hence, only accuracy metric is used as in [38].
TABLE II: The the generalization ability of the embedding weights learnt from STSB on SemAntoNeg Benchmark.
models acc $acc\_W_{stsb}$ gte-base-en-v1.5 47.12 47.17 (+0.05) gte-large-en-v1.5 53.58 56.18 (+2.60) bge-base-en-v1.5 53.67 56.78 (+3.11) bge-large-en-v1.5 64.68 68.73 (+4.05) e5-base-v2 54.60 56.41 (+1.81) e5-large-v2 60.83 60.87 (+0.04) multilingual-e5 68.87 73.14 (+4.27) mxbai-embed-large-v1 65.80 68.91 (+3.11) Average (Universal) 58.64 61.02 (+2.38) all-mpnet-base-v2 31.32 34.48 (+3.16) all-roBERTa-large-v1 35.32 37.17 (+1.85) Average (Contextual) 33.32 35.83 (+2.51)
Compared to STSB data with simple negation, SemAntoNeg Benchmark is more difficult as it contains negation, antonym and double negation (negation plus antonym). In this section, we evaluate the generalization capability of the simple negation awareness embedding weights, learned from STSB, on the SemAntoNeg Benchmark. The results are presented in Table II. The overall performance of universal text embeddings (with the mean performance of 58.64%) are much better than that of contextual text embedding baselines (with the mean performance of 33.32%). The random guess performance on SemAntoNeg Benchmark is 33.33%, which indicates that contextual text embeddings is prone to predicting the sentence with the highest lexical overlap while ignoring the negation [38]. On average (across all text embedding models), the embedding weights learnt from STSB improves the performance of the original embedding by 2.41%. This suggests that the weights learnt from STSB are able to generalize to SemAntoNeg Benchmark. However, the performance gains for e5-large-v2 and gte-base-en-v1.5 are less noticeable. In terms of model parameter size, it is observed that larger universal text embeddings demonstrate superior comprehension of complex negation compared to their smaller counterparts.
V-C Generalization ability of proposed solution on SemAntoNeg Benchmark
TABLE III: The generalization ability of proposed solution on SemAntoNeg Benchmark: $acc\_W_{anto}^{K}$ means the performance of embedding weights learnt from K training samples (K = 200, 500, 100). The mean and standard deviations of accuracy evaluated over 10 runs are shown in the table.
models acc $acc\_W_{anto}^{200}$ $acc\_W_{anto}^{500}$ $acc\_W_{anto}^{1000}$ gte-base-en-v1.5 $47.96$ $65.72$ $67.25$ $67.13$$\pm 0.49$$\pm 2.85$$\pm 0.98$$\pm 0.6$ gte-large-en-v1.5 $54.16$ $65.72$ $68.62$ $69.21$$\pm 0.49$$\pm 5.92$$\pm 4.3$$\pm 3.35$ bge-base-en-v1.5 $53.93$ $71.32$ $73.09$ $73.52$$\pm 0.25$$\pm 1.58$$\pm 0.61$$\pm 0.45$ bge-large-en-v1.5 $65.35$ $79.92$ 80.97 $80.86$$\pm 0.41$$\pm 1.09$$\pm 0.57$$\pm 0.35$ e5-base-v2 $54.58$ $71.34$ $71.96$ $72.1$$\pm 0.35$$\pm 1.2$$\pm 1.03$$\pm 0.79$ e5-large-v2 $61.5$ $75.24$ $76.97$ $77.01$$\pm 0.57$$\pm 1.87$$\pm 1.09$$\pm 0.8$ multilingual-e5 69.77 $80.10$ $80.88$ $80.66$$\pm 0.54$$\pm 1.67$$\pm 0.92$$\pm 1.14$ mxbai-embed-large-v1 $66.47$ 80.60 $80.93$ 80.99 ±$\pm 0.42$$\pm 0.74$$\pm 0.41$$\pm 0.36$ Average (Universal) 59.22 73.74 (+14.52) 75.08 (+15.86) 75.18 (+15.96) all-mpnet-base-v2 $32.1$ $45.22$ $45.33$ $45.66$$\pm 0.52$$\pm 0.99$$\pm 0.52$$\pm 0.78$ all-roBERTa-large-v1 $35.54$ $45.0$ $46.83$ $47.3$$\pm 0.45$$\pm 0.97$$\pm 0.84$$\pm 1.14$ Average (Contextual) 33.82 45.11 (+11.29) 46.08 (+12.26) 46.48 (+12.66)
In this section, the objective is to study whether the proposed embedding re-weighting solution is able to learn from complex negation (negation + antonym) from SemAntoNeg Benchmark. Unlike STSB, SemAntoNeg Benchmark has no predefined train test split. Hence, a stratified repeated random sampling approach is used to achieve a robust estimate of the performance. The random splitting is repeated 10 times, with 32% as training data (1000x4 samples) and 68% as test data (2152x4 samples). To test the data efficiency of the proposed solution, we test 3 different training data sizes: 200, 500 and 1000. The corresponding test performances (mean and standard deviations of accuracy evaluated over 10 runs) are shown in Table III as $acc\_W_{anto}^{200}$, $acc\_W_{anto}^{500}$, and $acc\_W_{anto}^{1000}$ respectively.
Generally speaking, the proposed weighting method is able to improve the performance on complex negation tasks across all text embedding models with small size training data. When the training data size is only 200 ($acc\_W_{anto}^{200}$ in Table III), the universal text embeddings’ performance are improved by 14.52% on average, while the contextual text embeddings’ performance are improved by 11.29%. When increasing the training data size from 200 to 500, the average universal text embeddings’ performance are further improved slightly by 1.34% with reduced standard deviations, while the average contextual text embeddings’ performance are further improved by 0.97%. Wilcoxon-Holm post hoc test with Critical Differences (CD) is also done to have an overall statistical comparison as shown in Figure 4: the proposed solution with three training sizes are all significantly better than the original embeddings, but there is no significant difference between the performance on 500 training datasize and 1000 training datasize (when alpha is 0.00001). Similar to the observations in previous sections: 1. larger universal text embeddings demonstrate superior comprehension of negation compared to their smaller counterparts, 2. universal embeddings benefit more from the proposed solution than contextual embeddings.
Figure 4: The CD diagram based on the Wilcoxon-Holm post hoc test result when alpha is 0.00001. Methods connected with bold black line have no significant difference.
V-D Generalization to Large Language Model based universal text embeddings
The previous sections have shown that the proposed dimension re-weighting method is able to efficiently improve the negation awareness of both universal and contextual text embeddings models. Compared to BERT-based universal embeddings tested in previous sections, LLM-based universal text embeddings use LLMs as backbones. These backbones are mostly decoder only models using causal attention mechanism which have larger parameter size (mostly 7B) [7]. The embedding dimensionality of most BERT-based universal and contextual embedding models is around 1000 while the embedding dimensionality of most LLM-based universal text embeddings is around 4000. However, LLM-based text embeddings have the advantage of good instruction following ability to deal with potentially mutually contradicted tasks and improve the generalization ability across different tasks, by adding a task specific instruction which describes the nature of the task (e.g. search relevant passages for the query) to the query side. In this case, the same model can provide different task-specific embeddings for different tasks (similar to multilingual-e5) [7].
The objective in this section is to study whether LLM-based high dimensional task-specific embeddings can be further improved by the proposed dimension weighting method. Three state-of-the-art LLM-based text embeddings are selected: SFR-Embedding-2_R with Mistral-7B [47] as backbone, gte-Qwen2-7B-instruct with Qwen2-7B [48] as backbone and gte-Qwen2-1.5B-instruct with Qwen2-1.5B [48] as backbone. The experimental results are shown in Table IV: Similar to the observations from previous sections, the proposed data efficient solution can improve significantly (around 12% improvement on average) the negation awareness of LLM-based text embeddings with small training data size.
On average, LLM-based universal text embeddings have better negation awareness than BERT-based universal text embeddings with around 6% improvement on SemAntoNeg Benchmark. However, this is primarily due to the good performance of gte-Qwen2-7B-instruct, while gte-Qwen2-1.5B-instruct and SFR-Embedding-2_R are not better than the average performance of BERT-based universal text embeddings with smaller parameter sizes. Among 4 embeddings with task-specific instructions (3 LLM-based embeddings and multilingual-e5), only multilingual-e5 and gte-Qwen2-7B-instruct show superior comprehension of complex negation compared to their similar-size counterparts. This indicates that different instruction-based embedding models have different instruction-following and instruction-generalization abilities. On the other side, even though gte-Qwen2-7B-instruct has good performance on SemAntoNeg (81.78%), the proposed embedding re-weighting method is still able to improve its performance by around 6%.
TABLE IV: The performance of proposed solution with LLM-based embedding models on SemAntoNeg Benchmark: $acc\_W_{anto}^{K}$ means the performance of embedding weights learnt from K training samples (K = 200, 500, 100). The mean and standard deviations of accuracy evaluated over 10 runs are shown in the table.
models acc $acc\_W_{anto}^{200}$ $acc\_W_{anto}^{500}$ $acc\_W_{anto}^{1000}$ SFR-Embedding-2_R $58.52$ $71.52$ $71.32$ $71.32$$\pm 0.52$$\pm 1.7$$\pm 0.85$$\pm 1.06$ gte-Qwen2-7B-instruct 81.78 87.43 87.52 87.44 ±$\pm 0.51$$\pm 0.36$$\pm 0.38$$\pm 0.46$ gte-Qwen2-1.5B-instruct $55.03$ $72.0$ $72.24$ $72.26$$\pm 0.52$$\pm 0.85$$\pm 0.47$$\pm 0.47$ Average 65.11 76.98 (+11.87) 77.03 (+11.92) 77.01 (+11.90)
V-E Summary
In this section, STSB dataset with simple negation is used firstly to test the effectiveness of the proposed dimension re-weighting method. The proposed solution is able to improve the accuracy of negation awareness task of both BERT-based universal text embeddings (4.68% improvement) and contextual text embeddings (3.83% improvement). The weights learnt on STSB are subsequently applied to SemAntoNeg Benchmark with complex negation (negation + antonym). The experimental results show that the STSB learnt weights can also improve the complex negation awareness with over 2% improvement on BERT-based universal text embedding models. Furthermore, the proposed solution is used to learn weights from small training data on SemAntoNeg Benchmark (200, 500 and 1000): performances based on all three training datasizes are significantly better than the original embedding (with over 10% improvement) across all BERT-based embedding models, which shows that the proposed solution is robust and data efficient. Finally, we assess the method’s effectiveness on three LLM-based embeddings. The results demonstrate that the proposed method can also improve the negation-awareness of high dimensional task-specific LLM-based universal embeddings.
Sustainability: The environmental impact of AI solutions has attracted a lot of attention in recent years [49, 50]. Compared to existing solutions in the literature, the proposed solution is significantly more sustainable in terms of training datasize, training time and computational cost.
VI Conclusion
Recent literature has witnessed the fast progress and development of universal text embeddings. However, due to the bias in popular evaluation benchmarks, the negation awareness capacity of these models remains unclear. In this work, we start with a holistic analysis in order to answer how similar are negated text pairs interpreted by the state-of-the-art universal text embeddings. The experimental results show that the state-of-the-art universal text embeddings lack negation awareness and interpret negated text pairs as roughly similar in terms of semantic meaning. To efficiently deal with the conflict that different tasks need different trade-off between topic and negation information among other semantic information, a data-efficient and sustainable embedding re-weighting function is proposed without modifying the parameters of text embedding models. The proposed solution is able to improve text embedding models’ negation awareness significantly on both STSB with simple negations and SemAntoNeg with complex negations. Furthermore, the proposed solution can also improve the negation awareness of LLM-based task-specific high dimensional embeddings. An additional benefit of the proposed solution is its flexibility: re-weighting is applied exclusively to tasks that need strong negation awareness, while other tasks like topic clustering/classification use the original unweighted embedding, which ensures the universality of universal text embeddings across different tasks. Future work will focus on generalizing the proposed solution to other NLP tasks requiring trade-offs among different types of semantic information. Another research direction is to combine the proposed data efficient solution with automatic training data generation by LLMs in order to improve the generalization ability of universal text embedding models on unseen tasks.